# Representation Learning with Conditional Information Flow Maximization

### v1

@@ B.1 Downstream Datasets and Tasks
The descriptions of 10 classification benchmarks are listed as follows: EmojiEval Barbieri et al. (2018) is designed for emoji prediction, which aims to predict its most likely emoji given a tweet. EmotionEval Mohammad et al. (2018) involves detecting the emotion evoked by a tweet and is based on the Affects in Tweets conducted during SemEval-2018. Following Barbieri et al. (2020), the most common four emotions (i.e., anger, joy, sadness, and optimism) are selected as the label sets. HatEval Basile et al. (2019) stems from SemEval-2019 HatEval and is used to predict whether a tweet is hateful towards immigrants or women. IronyEval Hee et al. (2018) is from SemEval-2018 Irony Detection and consists of identifying whether a tweet includes ironic intents or not. OffensEval Zampieri et al. (2019) is from SemEval-2019 OffensEval and involves predicting if a tweet contains any form of offensive language. SentiEval Rosenthal et al. (2017) comes from SemEval-2017 and includes data from previous runs (2013, 2014, 2015, and 2016) of the same SemEval task. The goal is to determine if a tweet is positive, negative, or neutral. StanceEval Mohammad et al. (2016) involves determining if the author’s text has a favorable, neutral, or negative position towards a proposition or target. ISEAR Scherer and Wallbott (1994) is from International Survey On Emotion Antecedents And Reactions project and contains reports on seven emotions each by close to 3000 respondents in 37 countries on all 5 continents. It aims to predict the emotion reaction. Due to the lack of a predefined split in the original corpus, we use the processed version by Hu et al. (2024) that randomly splits the dataset into train/valid/test sets in a ratio of 4:1:5 based on the label distribution. MELD Poria et al. (2019) contains multi-party conversation videos collected from Friends TV series. It is used to detect emotions in each utterance. [^5] MELD contains many types of context, including dialogue, speaker, and multi-modal signals. Different from other task-oriented methods Hu et al. (2021), this paper only considers the context-free textual utterance to better evaluate sentence classification performance. GoEmotions Demszky et al. (2020) is a corpus of comments from Reddit, with human annotations to 27 emotion categories or neutral. It is used fine-grained emotion prediction. Following Hu et al. (2024), nearly 16% of multi-label data was removed from the source corpus to better evaluate the performance of multi-class classification.
@@
Table 8: Ablation results (%) on 10 classification benchmarks. We bolded best results for each task.
@@
Table 10: Transferability results (%) of SentiEval $\rightarrow$ Others with RoBERTa backbone. Pre-train Obj. indicates the pre-training objective. Cls. refers to the classifier when fine-tuning.
@@ B.2 Comparison Methods
SVM Cortes and Vapnik (1995) is a machine learning algorithm with a hinge loss that aims to find the best hyperplane to separate data points into different classes. FastText Joulin et al. (2017) is an efficient classification method with negative log-likelihood loss based on n-gram features and a hierarchical softmax. BiLSTM is a bidirectional recurrent neural network Hochreiter and Schmidhuber (1997). GPT-3.5777https://chat.openai.com is an enhanced generative pre-trained transformer model based on text-davinci-003, optimized for chatting.

### v2

@@ B.1 Downstream Datasets and Tasks
The descriptions of 10 classification benchmarks are listed as follows: EmojiEval Barbieri et al. (2018) is designed for emoji prediction, which aims to predict its most likely emoji given a tweet. EmotionEval Mohammad et al. (2018) involves detecting the emotion evoked by a tweet and is based on the Affects in Tweets conducted during SemEval-2018. Following Barbieri et al. (2020), the most common four emotions (i.e., anger, joy, sadness, and optimism) are selected as the label sets. HatEval Basile et al. (2019) stems from SemEval-2019 HatEval and is used to predict whether a tweet is hateful towards immigrants or women. IronyEval Hee et al. (2018) is from SemEval-2018 Irony Detection and consists of identifying whether a tweet includes ironic intents or not. OffensEval Zampieri et al. (2019) is from SemEval-2019 OffensEval and involves predicting if a tweet contains any form of offensive language. SentiEval Rosenthal et al. (2017) comes from SemEval-2017 and includes data from previous runs (2013, 2014, 2015, and 2016) of the same SemEval task. The goal is to determine if a tweet is positive, negative, or neutral. StanceEval Mohammad et al. (2016) involves determining if the author’s text has a favorable, neutral, or negative position towards a proposition or target. ISEAR Scherer and Wallbott (1994) is from International Survey On Emotion Antecedents And Reactions project and contains reports on seven emotions each by close to 3000 respondents in 37 countries on all 5 continents. It aims to predict the emotion reaction. Due to the lack of a predefined split in the original corpus, we use the processed version by Hu et al. (2024) that randomly splits the dataset into train/valid/test sets in a ratio of 4:1:5 based on the label distribution. MELD Poria et al. (2019) contains multi-party conversation videos collected from Friends TV series. It is used to detect emotions in each utterance. The corpus contains many types of context, including dialogue, speaker, and multi-modal signals. Following Hu et al. (2024), this paper only considers the context-free textual utterance to better evaluate sentence classification performance. GoEmotions Demszky et al. (2020) is a corpus of comments from Reddit, with human annotations to 27 emotion categories or neutral. It is used fine-grained emotion prediction. Following Hu et al. (2024), nearly 16% of multi-label data was removed from the source corpus to better evaluate the performance of multi-class classification.
@@
Table 8: Ablation results (%) on 10 classification benchmarks. We bolded best results for each task under BERT and RoBERTa backbones.
@@
Table 10: Transferability results (%) of SentiEval $\rightarrow$ Others with RoBERTa backbone. Pre-train Obj. indicates the pre-training objective. Cls. refers to the classifier when fine-tuning. Best results for linear (i.e., Linear) and nonlinear (i.e., CNN) classifiers are highlighted in bold.
@@ B.2 Comparison Methods
SVM Cortes and Vapnik (1995) is a machine learning algorithm with a hinge loss that aims to find the best hyperplane to separate data points into different classes. FastText Joulin et al. (2017) is an efficient classification method with negative log-likelihood loss based on n-gram features and a hierarchical softmax. BiLSTM is a bidirectional recurrent neural network Hochreiter and Schmidhuber (1997). GPT-3.5666https://chat.openai.com is an enhanced generative pre-trained transformer model based on text-davinci-003, optimized for chatting.